A block of 100N is slowly slid up on a smooth incline of inclination 37 degree by a person .calculate the work done by the person in moving the block through a dostqnce of 2m,if the driving force is (a)parallel to the incline (b) in the horizontal direction.

Grade:11

1 Answers

Deepak
6 years ago
Firstly we resolve the components of weight along the inclined and perpendicular to the inclined. (a) parallel to the inclined component will be mgsin37°=100×(3÷5) {sin37°=3÷5} =60 work done=force ×displacement =60×2=120j similarly in horizontal direction mgcos37°=80{cos37°=4÷5} work done =force ×displacement =80×2=160j